the sum of two consecutive even integers is 34 less than their product. find the integers
Answer by ewatrrr(24785)

 
Hi,
two consecutive even integers: x and (x+2)
Question States***
2x + 2 = x(2+x)- 34
2x + 2 =2x + x^2 - 34
36 = x^2
x = ± 6
The two consecutive even integers are: 6, 8 0r -6, -4
CHECKING our answers***
14 = 48-34 and -10 = 24 - 34
Wish You the Best in your Studies.